A car radiator contains 30 quarts of 50% anti freeze. how many quarts should be drained and replaced with water so that the new solution is 30% antifreeze
Answer by ankor@dixie-net.com(22740)

A car radiator contains 30 quarts of 50% anti freeze.
how many quarts should be drained and replaced with water so that the new solution is 30% antifreeze
:
Let x = amt to be drained and amt of water to be added
:
.50(30-x) = .30(30)
15 - .5x = 9
15 - 9 = .5x
6 = .5x
x = 6/.5
x = 12 quarts removed and 12 qts of water added
:
See if that checks out
.50(30-12) = .30(30)
.5(18) = 9